Tags:
create new tag
view all tags

Calibration Tone Signal Stability

The variation in the calibration tone amplitude along the beamline is plotted below. A 3σ cut has been applied

calvar.png

Variation in the calibration tone phase:

calang.png

There seems to be a few BPMs for which, the calibration tone is not working stably. These are QM16FF, QM11FF and FB2FF. Their magnitudes (left) and phases (right) for the last 24 hours are shown below.

qm16ffcalm.png qm16ffcala.png

qm11ffcalm.png qm11ffcala.png

FB2FF caltone is at low amplitude and has large variation. It is also flipping in phase between ±π.

fb2ffcalm.png fb2ffcala.png

Plots of QM16FF and QM11FF on the same plot. There is no correlation, just coincidences. x and y seem to match OK.

fourcaltones.png fourcalangs.png

Results after tunnel access (10:15am) show QM16FF may have stabilised. Fingers crossed for QM11FF.

calmag20.png calang20.png

Caltones have been stable all weekend.

Friday/Saturday:

stabmsat.png stabasat.png

Sunday:

stabmsun.png stabasun.png

Edit | Attach | Watch | Print version | History: r4 < r3 < r2 < r1 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r4 - 23 Oct 2011 - FrankieCullinan

 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2022 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback